OpenStudy (anonymous):

30, 45, 68, 45, 58, 14 What is the mean?

OpenStudy (anonymous):

Add em up. Divide the answer by 6 To calculate an average of anything, just add up everything and divide by the number of numbers you have

OpenStudy (solomonzelman):

the mean is the average. for example if I had the following 4 numbers 3, 6, 11, 19 and wanted to find the mean, I would add these numbers up and divide by the "number of numbers" in this example it's 4. (3+ 6 + 11 + 20) / 4 = 40 / 4 = 10 So the mean would be 10. Can you do your's? I'll start you off. (30 + 45 + 68 + 45 + 58 + 14) / 6 = take it from here, and let me know if you need more help with this problem